The finance department at Victory Vacations has proposed a new search engine for use by their travel consultants. They charge their clients a flat $45 for each vacation consultation and pay their clerks $25.25 per hour.

Currently, a typical search process requires 60 minutes of each person's time. With the new tool, they will need only 40 minutes per search but they will be charged $2.90 per search in fees.

What is the percent change in multifactor productivity compared to the current search process? (Keep two decimal places)
Answer #1

Current productivity = 45 -25.25 / 60 min = 19.75 /hour

Proposed productivity =

40 min per search will give them $ [45 -2.9 -(25.25 x40/60)] = $25.26

hourly productivity = 25.26 x60/40 = 37.9

% increase in the MFP = (37.9/19.75 -1) x100 = 91.89%